In a cross between two heterozygotes, what is the probability that there will be two simultaneous offspring and one is homozygous recessive and the other homozygous dominant?

(hint: two events occuring at the same time)
Answer #1

Genotype of heterozygotes = Aa

Aa × Aa

Gametes A a
A AA homozygous dominant Aa heterozygous
a Aa heterozygous aa homozygous recessive

Probability of events occurring at same time = product rule

Probability of homozygous recessive = 1/4

Probability of homozygous dominant = 1/4

Probability of both of them occuring together = (1/4) × (1/4) = 1/16

Please rate.